OpenStudy (donnie1999):

What is the inverse of the function below? f(x) = 5x - 2

OpenStudy (anonymous):

Let's change f(x) to y so the equation is y=5x-2

OpenStudy (anonymous):

The inverse is when you switch y and x so that is x=5y-2

OpenStudy (anonymous):

Try and solve it should be easier

OpenStudy (donnie1999):

so would it be f-^1(x)=-2+5x ????

OpenStudy (anonymous):

No because when you divide the 5 you divide the whole equation

OpenStudy (anonymous):

so it's f^-1=(x-2)/5

OpenStudy (donnie1999):

SO IS IT f-^1(x)=x+2/5??

OpenStudy (donnie1999):

ohh srry i did not see u type that

OpenStudy (anonymous):

its alright yeah but keep in mind that both x and 2 are being divided by 5